Two radioactive sources A and B initially contian equal nno. of radioactive atoms. Source A has t½ of 1 hour and source B has t½ 2 hours. What is the ratio of rate of disintegration of A to B


i)initially


ii)after 2 hours


iii)after 10 hours


What is the method to do these questions?????

Grade:11

1 Answers

Srinivas Rao
13 years ago

Radio active decay follows first order kinetics.

So, rate of decay is given by Rate = k [N0 (1/2/)n]  where k is decay constant. k =0.693/ t1/2 ; n is number of half-lives.

(i) Initially, n=0

So, Rate -1 = (0.693/ 1) N0  and  Rate-2 = (0.693/ 2) N0 

Or (Rate-1 / Rate-2 ) = 2

(ii) After 2 Hours

2 Hours is 2 half lives for for A and 1 Half life for B

Rate -1 = (0.693/ 1) N0 (1/2)2 and  Rate-2 = (0.693/ 2) N0 (1/2)1

(Rate-1 / Rate-2 ) = 1

(iii) After 10 Hours

10 Hours is 10 half lives for for A and 5Half life for B

Rate -1 = (0.693/ 1) N0 (1/2)10 and  Rate-2 = (0.693/ 2) N0 (1/2)5

(Rate-1 / Rate-2 ) = 0.0625